If o(E) = 5:7, find o(E').
o(E') =
1
Expert's answer
2012-02-24T08:59:39-0500
E’ is an opposite set(event/probability) to the set(event/probability) E
So the sum of o(E)+ o(E') always equals 1
o(E)+ o(E') =1
5/7 + o(E') =1
o(E') =2/7=2:7